OpenStudy (anonymous):

If the integer m is divided by 6, the remainder is 5. What is the remainder if 4m is divided by 6?

OpenStudy (anonymous):

24 or 20 i believe

OpenStudy (anonymous):

Answer choices: A)0 B)1 C)2 D)4 E)5

OpenStudy (rc94):

Think of m as 6x + 5 where x can be any integer. This means that 4m = 4*(6x+5)=24x+20 24x+20=6(4x+3) + 2 so it will have remainder 2

OpenStudy (anonymous):

You could just pick numbers for this to make it easier. Since m%6 == 5 then you can say that m=11. With this you figure out 44%6 which is equal to 2

OpenStudy (anonymous):

^or that which is a better but harder way to solve it :p

OpenStudy (anonymous):

Thanks, you are truly a genius.